22% x (T-6660) = 19% x T
Need a value for T

Answer by 2897696(96)

22%(T-6660)=19%T
percent to decimal form, move decemial 2 places to the left:
22% =>0.22
19% =>0.19
now equation looks like:
0.22(T-6660)=0.19T
0.22T-1465.2=0.19T
-1465.2=-0.03T
48840=T
T=48840